# Do Teachers Dream of GenAI Widening Educational (In)equality? Envisioning the Future of K-12 GenAI Education from Global Teachers’ Perspectives

### v2

@@ Abstract.
Artificial Intelligence in Education, K-12 Education, Inequality
@@ 1. Introduction
Today, generative artificial intelligence (GenAI) is rapidly entering K–12 classrooms across the globe (Mintz et al., 2023; Mathur, 2025). Whether teachers welcome it or not, students are already engaging with GenAI tools in and beyond the classroom (Adisa and Adefisayo, 2025), for instance, using ChatGPT to draft essays (Levine et al., 2025), solve math problems (Walkington et al., 2025), or generate summaries of readings (Steele, 2023). This rapid spread makes GenAI not only a technological trend but also an urgent equity issue at the societal level. On one hand, it has the potential to democratize learning opportunities, providing personalized tutoring and cognitive support that could help close long-standing gaps in access and achievement (Tafazoli, 2024; James and Andrews, 2024). On the other hand, it risks exacerbating divides by disproportionately benefiting students with reliable infrastructure, digital literacy, and supportive school resources, while leaving behind those in under-resourced contexts (Hendawy, 2024; Murgia et al., 2024).
@@ 2.2. Tracing Educational (In)equalities: From Traditional Schooling to the GenAI Era
Prior HCI/CSCW work has shown that educational technologies do not translate into equitable outcomes automatically, but must be interpreted, negotiated, and appropriated by human actors within institutional and cultural constraints (Cao et al., 2025a; Orlikowski, 1992; Ackerman, 2000). Beyond empirical studies, policy and industry analyses such as Luckin and Holmes (2016)’s report in 2016, highlighting that AI in education will only benefit learners equitably when teachers are supported as orchestrators of AI use (Luckin and Holmes, 2016). Extending this perspective to nowadays, GenAI-in-school contexts, we aim to understand how teachers mobilize GenAI education to promote equity, and what organizational and societal conditions enable or hinder such work, connecting GenAI education to broader CSCW debates on infrastructure, labor, and mediation, and foregrounds teachers’ work as a gateway to designing equitable futures with AI.
@@ 3.1. Context: AI Education in the United States, South Africa, and Taiwan
South Africa’s AI education is shaped by stark socio-economic inequality, youth unemployment, and linguistic diversity (Spaull and Jansen, 2019; Statistics South Africa, 2025). Limited infrastructure and uneven school resources perpetuate inequalities in the distribution of national educational resources, especially between urban schools and township or rural schools (Sibuyi et al., 2024). To promote educational equity in the GenAI era and prepare the next generation for digital futures, government initiatives such as the Basic Education Employment Initiative (BEEI) have been introduced to improve digital readiness (Alexander, 2025). Teachers face the challenge of adapting AI use across 12 official languages while addressing unequal access to devices and training (Alexander, 2025). Although demand for AI is high (Mayowa Kuyoro and Umar Bagus and Anass Bensrhir and Ziyaad Bobat, 2025), local languages remain low-resource, and most tools in use are U.S.-developed (BusinessTech, 2024), which can limit their effectiveness in South Africa due to mismatches with local languages, curricula, culture and classroom contexts. Domestically, there are startups specialize in AI for education, but their growth is constrained by weak infrastructure, skills shortages, and regulatory uncertainty (Tracxn, 2025). In curriculum, IT is not compulsory: Grades R–9 include only basic digital literacy (Department of Education, South Africa, 2006), while Grades 10–12 offer two electives, Computer Applications Technology (CAT) and Information Technology (IT), with access highly unequal between elite and under-resourced schools (Department of Education, South Africa, 2006; Department of Basic Education, South Africa, 2015).
@@ 3.4. Data Analysis
We analyzed the interview data using a reflexive thematic analysis approach (Braun and Clarke, 2006, 2019). First, two members of the research team independently reviewed the transcripts and developed initial codes that captured recurring practices, challenges, and aspirations described by participants. We then iteratively refined the coding framework, combining inductive codes that emerged from the data with deductive codes derived from our research questions and prior work on AI education and educational inequality. To ensure reliability and depth, the team engaged in multiple rounds of coding comparison and discussion, resolving discrepancies through deliberation rather than statistical measures of inter-rater agreement, in line with traditional practices in qualitative HCI research (Xiao et al., 2025a). Themes were developed by clustering related codes, with particular attention to how teachers’ narratives reflected both global commonalities and context-specific conditions. Throughout the process, we wrote analytic memos to trace connections across regions and to foreground teachers’ own framings of how GenAI can either mitigate or exacerbate inequalities in K–12 education.
@@ 4. Findings: Equality-Oriented Teaching Practices in K–12 GenAI Education (RQ1)
Across these varied approaches, teachers framed their instructions around two interconnected goals related to educational (in)equality. First, they taught students to use GenAI to mitigate existing inequalities (e.g., accessing GenAI-generated materials when textbooks or other resources were scarce) so that disadvantaged learners would not fall further behind (see subsection 4.1). Second, they aimed to prevent GenAI from creating new inequalities by teaching responsible use, critical evaluation, cultural sensitivity, and the importance of maintaining human connection (see subsection 4.2). These dual orientations show how teachers viewed GenAI as a double-edged sword. On the one hand, it could open new opportunities for students who are often excluded from advanced technologies. On the other hand, if introduced without care, it could reinforce or even widen divides (Table 2).
@@ 7. Discussion
In the following discussion section, first, we reconceptualize GenAI education as a teacher practice of negotiating equality (see subsection 7.1). Second, we distinguish GenAI education from earlier forms of computing education, highlighting how its natural language interface lowers barriers for diverse learners but also introduces new equality challenges that demand teacher mediation (see subsection 7.2). Third, we extend beyond prior studies on individual adoption to demonstrate that teachers’ equality-oriented practices are deeply constrained by structural conditions of infrastructure, training, and social norms, underscoring the need to bridge micro-level pedagogical work with macro-level systemic inequalities (subsection 7.3). Finally, we contribute design and policy implications by outlining how GenAI systems and educational policies can better support teachers’ equality work, moving from individual initiative to systemic scaffolding (see section 8).
@@ 7.1. GenAI Education as a Teacher Practice of Negotiating Educational Equality
Second, our conceptualization positions teachers as central agents of sociotechnical negotiation. While prior HCI and learning sciences literature emphasizes systemic reforms, policy initiatives, or product design as levers of educational equality (Hutson et al., 2022; Uttamchandani, 2018), our study shows that equality is also forged through the micro-practices of teachers: the activities they design, the norms they reinforce, and the cultural framings they make explicit in their classrooms. Rather than viewing teachers as passive implementers of externally designed reforms, we argue that they act as mediators of technological possibility and as guardians of educational values, deciding when GenAI functions as an equalizer and when its risks must be curtailed. This builds on critical perspectives in CSCW and HCI that foreground human actors’ interpretive work in sociotechnical systems (Xiao et al., 2025b; Cheon, 2023), extending them to the context of K–12 education by showing how teachers’ local negotiations shape the equality trajectory of AI tools.
@@ 7.3. Structural Constraints on Equality-Oriented GenAI Education
8. Design and Policy Implications
Our study of equality-oriented GenAI education reveals that the pursuit of inclusivity is not merely a matter of classroom practice but requires supportive interventions in the design of GenAI tools and the policies that shape their adoption. Based on our findings, we identify several implications for both design and policy.
Our findings suggest that GenAI systems must be designed with equality as a central value rather than an afterthought, as teachers repeatedly pointed out that biased outputs and inaccessible interfaces constrained their ability to provide inclusive learning experiences. Such negative experiences, on the one hand, illustrate where current systems violate key Human–AI Interaction guidelines, pointing to design issues that GenAI product designers must address (Amershi et al., 2019). More specifically, these pitfalls inform us that design cannot stop at technical efficiency or personalization but must instead address accessibility, affordability, and cultural relevance. Affordable education-oriented versions should allow meaningful experimentation and iteration rather than restricting disadvantaged learners to minimal use. Likewise, GenAI models should incorporate diverse training data and localization features that reflect local languages, cultural references, and marginalized identities, ensuring that learners in Taiwan, South Africa, or elsewhere do not feel erased by homogenizing defaults. This mirrors postcolonial HCI arguments that technologies travelling across contexts must be re-designed with local infrastructures, languages, and epistemologies in mind to avoid reinforcing global inequities (Irani et al., 2010). Finally, design must move closer to pedagogy by embedding features that scaffold verification, critical reflection, and responsible classroom use, helping teachers teach not only how to operate GenAI but how to use it fairly and critically. Moreover, teachers’ negative experiences highlight the need for stronger AI literacy among developers, managers, and decision-makers so that potential harms can be anticipated rather than treated as afterthoughts, as emphasized by CHI researchers (Xie et al., 2025).
At the same time, teachers’ efforts were deeply circumscribed by structural conditions that cannot be resolved within the classroom alone, highlighting the need for policy interventions at government, community, and school levels. Infrastructural investment is needed to ensure that device distribution is paired with reliable connectivity, particularly in rural and low-income communities-a persistent challenge highlighted in CSCW research on infrastructural inequality, especially in cultural or geographic minority contexts (Sultana et al., 2019; Star, 1999) and cannot be overlooked in future research. Teacher capacity likewise requires systemic support, with professional development that goes beyond tool introduction toward sustained pedagogical frameworks for responsible GenAI integration (Ding et al., 2024). Our findings also underscore the importance of community trust: parental skepticism, cultural taboos, and political resistance shaped students’ access to GenAI, suggesting the need for public education and participatory design, consistent with CSCW work involving community-centered design (Villari and others, 2021; Zytko et al., 2022). Finally, equity can to be addressed across schools with varied levels of resources, not only within them, according to interviewees’ success stories. Thus sharing resources and capacity-building to avoid the Matthew Effect where privileged schools accelerate while others fall behind.
9. Limitations and Future Works

### v3

@@ Abstract.
Artificial Intelligence in Education, Education Equality, K-12 Education, Global Education, Inequality, Teacher, Generative AI
@@ 1. Introduction
Today, generative artificial intelligence (GenAI) is rapidly entering K–12 classrooms across the globe (Mintz et al., 2023; Mathur, 2025). Whether teachers welcome it or not, students are already engaging with GenAI tools in and beyond the classroom (Adisa and Adefisayo, 2025), for instance, using ChatGPT to draft essays (Levine et al., 2025), solve math problems (Walkington et al., 2025), or generate summaries of readings (Steele, 2023). This rapid spread makes GenAI not only a technological trend but also an urgent equality issue at the societal level. On one hand, it has the potential to democratize learning opportunities, providing personalized tutoring and cognitive support that could help close long-standing gaps in access and achievement (Tafazoli, 2024; James and Andrews, 2024). On the other hand, it risks exacerbating divides by disproportionately benefiting students with reliable infrastructure, digital literacy, and supportive school resources, while leaving behind those in under-resourced contexts (Hendawy, 2024; Murgia et al., 2024).
@@ 2.2. Tracing Educational (In)equalities: From Traditional Schooling to the GenAI Era
Prior HCI/CSCW works have shown that educational technologies do not translate into equitable outcomes automatically, but must be interpreted, negotiated, and appropriated by human actors within institutional and cultural constraints (Cao et al., 2025a; Orlikowski, 1992; Ackerman, 2000). Beyond empirical studies, policy and industry analyses such as Luckin and Holmes (2016)’s report in 2016, highlighting that AI in education will only benefit learners equitably when teachers are supported as orchestrators of AI use (Luckin and Holmes, 2016). Extending this perspective to nowadays, GenAI-in-school contexts, we aim to understand how teachers mobilize GenAI education to promote equality, and what organizational and societal conditions enable or hinder such work, connecting GenAI education to broader CSCW debates on infrastructure, labor, and mediation, and foregrounds teachers’ work as a gateway to designing equitable futures with AI.
@@ 3.1. Context: AI Education in the United States, South Africa, and Taiwan
South Africa’s AI education is shaped by stark socio-economic inequality, youth unemployment, and linguistic diversity (Spaull and Jansen, 2019; Statistics South Africa, 2025). Limited infrastructure and uneven school resources perpetuate inequalities in the distribution of national educational resources, especially between urban schools and township or rural schools (Sibuyi et al., 2024). To promote educational equality in the GenAI era and prepare the next generation for digital futures, government initiatives such as the Basic Education Employment Initiative (BEEI) have been introduced to improve digital readiness (Alexander, 2025). Teachers face the challenge of adapting AI use across 12 official languages while addressing unequal access to devices and training (Alexander, 2025). Although demand for AI is high (Mayowa Kuyoro and Umar Bagus and Anass Bensrhir and Ziyaad Bobat, 2025), local languages remain low-resource, and most tools in use are U.S.-developed (BusinessTech, 2024), which can limit their effectiveness in South Africa due to mismatches with local languages, curricula, culture and classroom contexts. Domestically, there are startups specialize in AI for education, but their growth is constrained by weak infrastructure, skills shortages, and regulatory uncertainty (Tracxn, 2025). In curriculum, IT is not compulsory: Grades R–9 include only basic digital literacy (Department of Education, South Africa, 2006), while Grades 10–12 offer two electives, Computer Applications Technology (CAT) and Information Technology (IT), with access highly unequal between elite and under-resourced schools (Department of Education, South Africa, 2006; Department of Basic Education, South Africa, 2015).
@@ 3.4. Data Analysis
We analyzed the interview data using a reflexive thematic analysis approach (Braun and Clarke, 2006, 2019). First, two members of the research team independently reviewed the transcripts and developed initial codes that captured recurring practices, challenges, and aspirations described by participants. We then iteratively refined the coding framework, combining inductive codes that emerged from the data with deductive codes derived from our research questions and prior work on AI education and educational inequality. To ensure reliability and depth, the team engaged in multiple rounds of coding comparison and discussion, resolving discrepancies through deliberation rather than statistical measures of inter-rater agreement, in line with traditional practices in qualitative HCI research (Xiao et al., 2025b). Themes were developed by clustering related codes, with particular attention to how teachers’ narratives reflected both global commonalities and context-specific conditions. Throughout the process, we wrote analytic memos to trace connections across regions and to foreground teachers’ own framings of how GenAI can either mitigate or exacerbate inequalities in K–12 education.
@@ 3.5. Positionality
3.6. Ethical and Reflexive Considerations
In addition to formal IRB approval, this study required ongoing ethical reflection due to the sensitive nature of K–12 education and the involvement of teachers working with legally and socially protected populations. While our interviews focused on teachers’ pedagogical practices rather than collecting data from students directly, teachers’ accounts frequently implicated issues of student data, surveillance, and emotional vulnerability in classroom uses of GenAI. We therefore approached data collection and analysis with heightened attention to privacy, power, and responsibility.
First, data privacy emerged as an implicit but significant ethical concern. Teachers often described using commercial GenAI platforms whose data practices were opaque to them, particularly with respect to how student prompts, outputs, or interaction logs might be stored or reused. Although our study did not examine platform data flows directly, we recognize that teachers’ uncertainty around data governance shaped both their teaching practices and their ethical hesitation. In reporting these accounts, we deliberately avoided collecting or reproducing any identifiable student data and treated teachers’ descriptions of classroom incidents at an abstract level to prevent indirect disclosure of sensitive information.
Second, our cross-national design foregrounded power asymmetries that operate differently across policy regimes. Teachers in relatively centralized systems (e.g., Taiwan) and decentralized systems (e.g., the United States) described distinct forms of constraint, yet across contexts, they shared a common position of being held responsible for ethical AI use without having commensurate authority over infrastructural decisions, platform selection, or data policies. As researchers, we are attentive to how documenting teachers’ struggles with GenAI risks normalizing this unequal distribution of responsibility. To mitigate this, our analysis consistently situates teachers’ practices within broader institutional and policy contexts rather than framing ethical challenges as individual shortcomings.
Third, we reflect on our positionality as researchers studying ethically contested technologies. While teachers often asked normative questions about what constitutes “responsible” GenAI use in classrooms, our role was not to prescribe correct practices but to surface the tensions and constraints that shape teachers’ decision-making.
Together, these ethical and reflexive considerations reinforce one of our central claims: that equality-oriented GenAI education is shaped not only by pedagogical intentions, but also by asymmetric power relations and unresolved governance questions surrounding data, accountability, and control. Making these dynamics explicit strengthens the interpretive validity of our findings and underscores the need for ethical responsibility to be addressed at systemic, rather than solely individual, levels.
@@ 4. Findings: Equality-Oriented Teaching Practices in K–12 GenAI Education (RQ1)
Across these varied approaches, teachers framed their instructions around two interconnected goals related to educational (in)equality. First, they taught students to use GenAI to mitigate existing inequalities (e.g., accessing GenAI generated materials when textbooks or other resources were scarce) so that disadvantaged learners would not fall further behind (see subsection 4.1). Second, they aimed to prevent GenAI from creating new inequalities by teaching responsible use, critical evaluation, cultural sensitivity, and the importance of maintaining human connection (see subsection 4.2). These dual orientations show how teachers viewed GenAI as a double-edged sword. On the one hand, it could open new opportunities for students who are often excluded from advanced technologies. On the other hand, if introduced without care, it could reinforce or even widen divides (Table 2).
@@ 7. Discussion
In the following discussion section, first, we reconceptualize GenAI education as a teacher practice of negotiating equality (see subsection 7.1). Second, we distinguish GenAI education from earlier forms of computing education, highlighting how its natural language interface lowers barriers for diverse learners but also introduces new equality challenges that demand teacher mediation (see subsection 7.2). Third, we extend beyond prior studies on individual adoption to demonstrate that teachers’ equality-oriented practices are deeply constrained by structural conditions of infrastructure, training, and social norms, underscoring the need to bridge micro-level pedagogical work with macro-level systemic inequalities (subsection 7.3). Finally, we synthesize these findings into a set of design tensions that articulate broader HCI contributions, and derive design and policy implications for supporting equality-oriented GenAI education at scale (subsection 7.4, subsection 7.5).
@@ 7.1. GenAI Education as a Teacher Practice of Negotiating Educational Equality
Second, our conceptualization positions teachers as central agents of sociotechnical negotiation. While prior HCI and learning sciences literature emphasizes systemic reforms, policy initiatives, or product design as levers of educational equality (Hutson et al., 2022; Uttamchandani, 2018), our study shows that equality is also forged through the micro-practices of teachers: the activities they design, the norms they reinforce, and the cultural framings they make explicit in their classrooms. Rather than viewing teachers as passive implementers of externally designed reforms, we argue that they act as mediators of technological possibility and as guardians of educational values, deciding when GenAI functions as an equalizer and when its risks must be curtailed. This builds on critical perspectives in CSCW and HCI that foreground human actors’ interpretive work in sociotechnical systems (Xiao et al., 2025c; Cheon, 2023), extending them to the context of K–12 education by showing how teachers’ local negotiations shape the equality trajectory of AI tools.
@@ 7.3. Structural Constraints on Equality-Oriented GenAI Education
These findings clarify why equality-oriented GenAI education cannot rely on teachers’ individual commitment alone. While teachers actively negotiate GenAI’s role in alleviating and preventing inequalities through everyday pedagogical practices, the scope and sustainability of this work are fundamentally shaped by forces beyond the classroom. The constraints identified in this study, ranging from infrastructural gaps and uneven training to contested social norms, systematically delimit what teachers can achieve through local negotiation. These structural conditions echo prior work showing that educational technologies do not automatically translate into equitable outcomes, but must be interpreted and appropriated within broader institutional and cultural contexts (Ackerman, 2000; Orlikowski, 1992; Cao et al., 2025a). This gap between teachers’ equality work and the structural conditions that enable or constrain it directly motivates our design tensions and implications below. Supporting equitable GenAI education therefore requires moving beyond individual initiative toward systemic scaffolding, in which GenAI systems are designed with equality as a core value and policies intervene to redistribute resources, build capacity, and cultivate public trust.
7.4. Design Tensions in Equality-Oriented GenAI Education
Synthesizing findings across RQ1–RQ3, we identify a set of design tensions that structure equality-oriented GenAI education. Making these tensions explicit helps explain why teachers’ practices are both creative and constrained, and why equality-oriented GenAI education cannot be reduced to teachers’ best practices alone.
7.4.1. Lowering barriers versus deepening new forms of inequality.
GenAI’s natural language interface lowers technical entry barriers and enables broader participation, particularly for students with limited digital literacy or learning disabilities (RQ1). However, the same accessibility also introduces new inequalities based on usage patterns, such as shortcut reliance, uneven critical engagement, and differential exposure to AI harms (RQ1). Teachers were therefore caught in a tension between promoting access and guarding against superficial or harmful use. This tension highlights that inclusion through accessibility does not automatically translate into equitable learning outcomes, and that GenAI systems designed to be “easy to use” still require pedagogical scaffolding to avoid reproducing stratified forms of participation.
From a broader HCI perspective, this tension challenges the common assumption that lowering interactional barriers is inherently equality-promoting (Kim et al., 2025; Xiao et al., 2025a; Harrington et al., 2019). Our findings suggest that accessibility-oriented design decisions can simultaneously redistribute participation and stratify outcomes, calling for HCI research to more closely examine how ease of use reshapes patterns of engagement, dependence, and harm across different user groups.
7.4.2. Personalization versus standardization of educational values.
Teachers leveraged GenAI to personalize learning for marginalized students, enabling flexible pacing, tailored explanations, and individualized planning (RQ1). At the same time, they encountered standardized outputs that reflected dominant cultural, linguistic, or epistemic norms, marginalizing local identities and minority experiences (RQ1, RQ2). This tension placed teachers in the role of cultural interpreters who had to correct, contextualize, or resist AI-generated representations. From a design perspective, this reveals a fundamental challenge: personalization at the level of individual learners can coexist with homogenization at the level of cultural representation, unless localization and cultural sensitivity are treated as core design commitments.
For HCI, this tension extends beyond education by foregrounding personalization as a multi-level phenomenon. While systems may successfully tailor content to individual users, they can simultaneously standardize values and representations at scale (Li et al., 2025; Xiao et al., 2025d; Schecter and Richardson, 2025). This insight urges HCI researchers to examine not only who personalization serves, but also whose cultural and epistemic assumptions are embedded and amplified through personalized AI systems.
7.4.3. Teacher mediation versus structural dependency.
Across contexts, teachers actively mediated GenAI use by setting norms, disciplining misuse, foregrounding bias, and emphasizing human care (RQ1). Yet their capacity to sustain this mediation was repeatedly constrained by infrastructural gaps, insufficient training, and restrictive social norms beyond their control (RQ2, RQ3). This tension exposes a misalignment between the moral and pedagogical responsibility placed on teachers and the limited structural support available to them. While GenAI education is often framed as a matter of teacher competence or innovation, our findings show that equality-oriented mediation depends on systemic conditions that cannot be addressed through individual effort alone.
This tension contributes a general insight for HCI and CSCW research on human-in-the-loop systems. It highlights a recurring pattern in which human actors are positioned as ethical and relational safeguards for AI systems, while the infrastructural and organizational conditions required to support this work remain underdeveloped. Rather than treating human mediation as an unqualified design asset, HCI must interrogate when such mediation becomes an uneven redistribution of accountability that obscures structural responsibility.
7.4.4. Human care versus machine-mediated support.
Finally, teachers grappled with the tension between GenAI as a supportive tool and its potential to displace human relationships, particularly in emotional and relational domains (RQ1). While some students used GenAI as a source of reassurance or guidance, teachers emphasized that emotional care, moral development, and social learning remain fundamentally human responsibilities. This tension underscores that GenAI education is not only about knowledge or skills, but also about preserving relational values that underpin equitable education. Designing GenAI systems for educational contexts therefore requires careful attention to where automation should stop and where human judgment and care must remain central.
More broadly, this tension informs HCI debates on automation boundaries in socially consequential domains. Our findings suggest that affective and relational dimensions of interaction are central to equality and inclusion when AI systems enter contexts involving guidance and moral development (Fan et al., 2025; Henriksen et al., 2025; Yu et al., 2025; Fan et al., 2024).
7.5. Design and Policy Implications
The design tensions identified above point to a central implication: equality-oriented GenAI education cannot be achieved through K-12 teachers’ classroom practice alone, nor through isolated improvements to AI systems. Instead, addressing these tensions requires coordinated interventions at both the design and policy levels. In this section, we translate the tensions between access and inequality, personalization and standardization, teacher mediation and structural dependency, and human care and machine support into concrete implications for GenAI system design and educational policy.
7.5.1. Implications for GenAI System Design.
Several of the tensions identified in Section 7.4—particularly those between lowering barriers and deepening new inequalities, personalization and standardization of values, and human care versus machine-mediated support—directly implicate the design of GenAI systems.
Our findings suggest that GenAI systems must be designed with equality as a central value rather than an afterthought, as teachers repeatedly pointed out that biased outputs and inaccessible interfaces constrained their ability to provide inclusive learning experiences. Such negative experiences illustrate where current systems violate key Human–AI Interaction guidelines, pointing to design issues that GenAI product designers must address (Amershi et al., 2019). More specifically, these pitfalls show that design cannot stop at technical efficiency or surface-level personalization, but must attend to how accessibility, affordability, and cultural relevance interact with educational values.
For example, affordable education-oriented versions of GenAI should allow meaningful experimentation and iteration, rather than restricting disadvantaged learners to minimal or heavily constrained use. Likewise, GenAI models should incorporate diverse training data and localization features that reflect local languages, cultural references, and marginalized identities, ensuring that learners in Taiwan, South Africa, or elsewhere do not feel erased by homogenizing defaults. This directly responds to the tension between personalization and standardization identified earlier, and echoes postcolonial HCI arguments that technologies traveling across contexts must be redesigned with local infrastructures, languages, and epistemologies in mind to avoid reinforcing global inequities (Irani et al., 2010; Ahmed et al., 2015; Ferreira, 2024).
Finally, addressing the tension between human care and machine-mediated support requires design to move closer to pedagogy. GenAI systems for educational contexts should embed features that scaffold verification, critical reflection, and responsible classroom use, helping teachers teach not only how to operate GenAI but how to engage with it ethically and critically (Xiao et al., 2025d). Teachers’ negative experiences also highlight the need for stronger AI literacy among developers, managers, and decision-makers, so that potential harms, especially those affecting vulnerable learners, can be anticipated rather than treated as afterthoughts, as emphasized by prior CHI research (Xie et al., 2025).
7.5.2. Implications for Educational Policy.
At the same time, the tension between teacher mediation and structural dependency makes clear that many constraints shaping equality-oriented GenAI education cannot be resolved through design choices or individual teacher effort alone. These challenges call for policy interventions at government, community, and school levels.
Infrastructural investment is needed to ensure that device distribution is paired with reliable connectivity, particularly in rural and low-income communities, a persistent challenge highlighted in HCI and CSCW research on infrastructural inequality, especially in cultural or geographic minority contexts (Sultana et al., 2019; Star, 1999). Without such investment, the promise of lowered interactional barriers risks reproducing new forms of exclusion.
Teacher capacity likewise requires systemic support. Professional development must go beyond tool introduction toward sustained pedagogical frameworks for responsible GenAI integration, enabling teachers to manage the tensions between access, critical engagement, and ethical use over time (Ding et al., 2024). Our findings also underscore the importance of community trust: parental skepticism, cultural taboos, and political resistance shaped students’ access to GenAI, suggesting the need for public education and participatory approaches to policy-making, consistent with HCI and CSCW work on community-centered design (Villari and others, 2021; Zytko et al., 2022).
Finally, equity must be addressed not only within individual schools but across schools with highly uneven resources. Interviewees’ success stories illustrate the importance of cross-school resource sharing and capacity building to avoid a Matthew Effect, in which already privileged schools accelerate their GenAI adoption while under-resourced schools fall further behind. Policy interventions that support collaboration, redistribution, and long-term capacity building are therefore essential for translating teachers’ equality-oriented visions into sustainable practice.
8. Limitations and Future Work